# Diego Marcondes 1122 $0$-0002-6087-4821    Mariana Feldman 11 0009-0002-1961-9172    Junior Barrera 11 $0$-0003-0439-0475

### v1

@@ Abstract
With the advent of deep learning, there have been attempts to insert mathematical morphology (MM) operators into convolutional neural networks (CNN), and the most successful endeavor to date has been the morphological neural networks (MNN). Although MNN have performed better than CNN in solving some problems, they inherit their black-box nature. Furthermore, in the case of binary images, they are approximations, which loose the Boolean lattice structure of MM operators and, thus, it is not possible to represent a specific class of W-operators with desired properties. In a recent work, we proposed the Discrete Morphological Neural Networks (DMNN) for binary image transformation to represent specific classes of W-operators and estimate them via machine learning. We also proposed a stochastic lattice gradient descent algorithm (SLGDA) to learn the parameters of Canonical Discrete Morphological Neural Networks (CDMNN), whose architecture is composed only of operators that can be decomposed as the supremum, infimum, and complement of erosions and dilations. In this paper, we propose an algorithm to learn unrestricted sequential DMNN (USDMNN), whose architecture is given by the composition of general W-operators. We consider the representation of a W-operator by its characteristic Boolean function, and then learn it via a SLGDA in the Boolean lattice of functions. Although both the CDMNN and USDMNN have the Boolean lattice structure, USDMNN are not as dependent on prior information about the problem at hand, and may be more suitable in instances in which the practitioner does not have strong domain knowledge. We illustrate the algorithm in a practical example.
@@
discrete morphological neural networks image processing mathematical morphology U-curve algorithms stochastic lattice gradient descent
@@ 1 Introduction
More recently, mathematical morphology methods have been studied in connection with deep neural networks, by either combining convolutional neural networks (CNN) with a morphological operator [10], or by replacing the convolution operations of CNN with basic morphological operators, such as erosions and dilations, obtaining the so-called morphological neural networks (MNN) [7, 9, 13]. A MNN has the general framework of neural networks, and its specificity is on the fact that the layers realize morphological operations. Although it has been seen empirically that convolutional layers could be replaced by morphological layers [7], and MNN have shown a better performance than CNN in some tasks [9], they are not more interpretable than an ordinary CNN. In the context of machine learning theory, the interpretability is related to the notion of hypothesis space, which is a key component of machine learning theory.
@@
In this context, in [11] we proposed the Discrete Morphological Neural Networks (DMNN) for binary image transformation to represent translation invariant and locally defined operators, i.e., W-operators, and estimate them via machine learning. A DMNN architecture is represented by a Morphological Computational Graph that combines operators via composition, infimum, and supremum to build more complex operators. In [11] we proposed a stochastic lattice gradient descent algorithm (SLGDA) to train the parameters of Canonical Discrete Morphological Neural Networks (CDMNN) based on a sample of input and output images under the usual machine learning approach. The architecture of a CDMNN is composed only of canonical operations (supremum, infimum, or complement) and operators that can be decomposed as the supremum, infimum, and complement of erosions and dilations with the same structural element. The DMNN is a true mathematical morphology method since it retains the control over the design and the interpretability of results of classical mathematical morphology methods, which is a relevant advantage over CNN.
@@
In this paper, we propose an algorithm to train unrestricted sequential DMNN (USDMNN), whose architecture is given by the composition of general W-operators. The algorithm considers the representation of a W-operator by its characteristic Boolean function, which is learned via a SLGDA in the Boolean lattice of functions. This SLGDA differs from that of [11] which minimizes an empirical error in a lattice of sets and intervals that are the structural elements and intervals representing operators such as erosions, dilations, openings, closings and sup-generating.
@@
As opposed to CDMNN, which can be designed from prior information to represent a constrained class of operators, USDMNN are subject to overfitting the data, since it can represent a great class of operators, that may fit the data, but not generalize to new examples. In order to address this issue, we control the complexity of the architecture by selecting from data the window of each W-operator in the sequence (layer). By restricting the window, we create equivalence classes on the characteristic function domain that constrain the class of operators that each layer can represent, decreasing the overall complexity of the operators represented by the architecture and mitigating the risk of overfitting. We propose a SLGDA to select the windows by minimizing a validation error in a lattice of sets within the usual model selection framework in machine learning.
@@
In Section 2, we present some notations and definitions, and in Section 3 we formally define the USDMNN that is a particular example of the DMNN proposed by [11]. In Section 4, we present the SLGDAs to learn the windows and the characteristic functions of the W-operators in a USDMNN, and in Section 5 we apply the USDMNN to the same dataset used in [11] in order to compare the results. In Section 6, we present our conclusions.
@@ 2 W-operators
Let $W$ be a finite subset of $E$. A set operator $\psi$ is called locally defined within a window $W$ if, and only if, $\forall h\in E$, $h\in\psi(X)\iff h\in\psi(X\cap W_{h})$ for $X\in\mathcal{P}(E)$. The collection $\Psi_{W}$ of t.i. operators locally defined within a window $W\in\mathcal{P}(E)$ inherits the complete lattice structure of $(\mathcal{P}(E),\subseteq)$ by setting, $\forall\psi_{1},\psi_{2}\in\Psi$, {linenomath}
@@
Define by $\Omega=\cup_{W\in\mathcal{P}(E),|W|<\infty}\Psi_{W}$ the subset of $\Psi$ of all operators from $\mathcal{P}(E)$ to $\mathcal{P}(E)$ that are t.i. and locally defined within some finite window $W\in\mathcal{P}(E)$. The elements of $\Omega$ are called $W$-operators.
@@
$T(\psi)(X)=\begin{cases}1,&\text{ if }o\in\psi(X)\\ 0,&\text{ otherwise}.\end{cases}$ $\psi\in\Psi_{W},X\in\mathcal{P}(W).$
@@ 3.1 Sequential Morphological Computational graph
=$\mathcal{E}=\left\{(\mathfrak{v}_{i},\mathfrak{v}_{1}),(\mathfrak{v}_{n},\mathfrak{v}_{o})\right\}\bigcup\left\{(\mathfrak{v}_{j},\mathfrak{v}_{j+1}):j=1,\dots,n-1\right\},$
$\mathcal{C}(\mathfrak{v}_{i})=\mathcal{C}(\mathfrak{v}_{o})=\iota$ and $\mathcal{C}(\mathfrak{v}_{j})\in\Omega,\ j=1,\dots,n$. This computational graph satisfies the axioms of MCG (see [11] for more details).
@@
Denote by $\psi_{\mathcal{G}}(X)$ the output of vertex $\mathfrak{v}_{o}$ when the input of vertex $\mathfrak{v}_{i}$ is $X\in\mathcal{P}(E)$ as the set operator $\psi_{\mathcal{G}}:\mathcal{P}(E)\to\mathcal{P}(E)$ generated by MCG $\mathcal{G}$. The operator $\psi_{\mathcal{G}}$ is actually t.i. and locally defined within a window $W_{\mathcal{G}}$ (cf. Proposition 5.1 in [11]). We define the sequential Discrete Morphological Neural Network represented by $\mathcal{G}$ as the translation invariant and locally defined set operator $\psi_{\mathcal{G}}$.
3.2 Sequential Discrete Morphological Neural Networks Architectures
@@ 3.3 Representation of USDMNN
in which $\psi_{j}=\mathcal{C}(\mathfrak{v}_{j}),j=1,\dots,n$. Based on (2), we propose a representation for the class of operators realized by an USDMNN.
@@
Since an operator locally defined in $W$ is also locally defined in any $W^{\prime}\supset W$ (cf. Proposition 5.1 in [3]) it follows that $\Theta$ is actually an overparametrization of $\mathcal{H}(\mathcal{A})$ since there are many representations $(W^{\prime},f_{\psi})$ for a same W-operator locally defined in $W\subsetneq F_{d_{i}}$. This overparametrization also happens in the representation of CDMNN proposed in [11] and we take advantage of it to propose a SLGDA to train USDMNN and mitigate the risk of overfitting.
4 Training USDMNN via the stochastic lattice gradient descent algorithm
@@
The U-curve algorithms [1, 6, 14, 15] were proposed to minimize a U-curve function on a Boolean lattice. In summary, these algorithms perform a greedy search of a Boolean lattice, at each step jumping to the point at distance one with the least value of the function, and stopping when all neighbor points have a function value greater or equal to that of the current point. This greedy search of a lattice, that at each step goes to the direction that minimizes the function, is analogous to the dynamic of the gradient descent algorithm to minimize a function with domain in $\mathbb{R}^{p}$. Inspired by the U-curve algorithms and by the success of stochastic gradient descent algorithms for minimizing overparametrized functions in $\mathbb{R}^{p}$, and following the ideas of [11], we propose a stochastic lattice gradient descent algorithm (SLGDA) to train USDMNN. In Figure 1, we present the main ideas of the algorithm, and in Sections 4.1 and 4.2 we formally define it.
Figure 1: Illustration of the deterministic version of the SLGDA to (a) learn the USDMNN windows and (b) train a USDMNN with fixed windows. To simplify the illustration, we considered only the possibility of adding a point to a window, and flipping a bit from 0 to 1 of a characteristic function, at each step, even though a point can be erased from a window, and a flipping from 1 to 0 may happen, if they have the least respective error. The training error $L_{t}$ or validation error $L_{v}$ of each point is on top of it.
@@ 4.1 Training a USDMNN with fixed windows
The SLGDA for learning the characteristic functions of a USDMNN with fixed windows is formalized in Algorithm 1. The initial point $\theta\in\Theta_{\boldsymbol{W}}$, a batch size $b$, the number $n$ of neighbors to be sampled at each step, and the number of training epochs are fixed. The initial point is stored as the point with minimum training error visited so far. For each epoch, the training sample is randomly partitioned in $N/b$ batches, and we denote the training error on batch $j$ by $L_{t}^{(j)}$. For each batch, $n$ neighbors of $\theta$ are sampled and $\theta$ is updated to the sampled neighbor with the least training error $L_{t}^{(j)}$, that is calculated on the sample batch $j$. Observe that $\theta$ is updated at each batch, so during an epoch, it is updated $N/b$ times. At the end of each epoch, the training error $L_{t}(\theta)$ of $\theta$ on the whole training sample is compared with the error of the point with the least training error visited so far at the end of an epoch, and it is stored as this point if its training error is lesser. After the predetermined number of epochs, the algorithm returns the point with the least training error on the whole sample visited at the end of an epoch.
Observe that Algorithm 1 has two sources of stochasticity: the sample of neighbors and the sample batches. If $n=n(\theta)\coloneqq|N(\theta)|$ and $b=N$, then this algorithm reduces to the deterministic one illustrated in Figure 1. Furthermore, the complexity of the algorithm is controlled by the number $n$ of sampled neighbors, the batch size $b$ and the number of epochs. See [11] for a further discussion about a SLGDA.
Algorithm 1 Stochastic lattice gradient descent algorithm for learning the characteristic functions of a USDMNN with fixed windows $\boldsymbol{W}=\{W_{1},\dots,W_{n}\}$.
@@
4.2 Learning the USDMNN windows
In order to learn the windows of a USDMNN, we apply the SLGDA on $\mathscr{C}$, that is a subset of the Boolean lattice $\prod_{i=1}^{n}\{0,1\}^{F_{d_{i}}}$. For each $\boldsymbol{W}\in\mathscr{C}$, let $L_{v}(\boldsymbol{W})\coloneqq L_{v}(\theta^{\mathbb{A}}_{\boldsymbol{W}})$ be the validation error of the USDMNN realized by $\theta^{\mathbb{A}}_{\boldsymbol{W}}$, which was learned by Algorithm 1.
@@ 4.2 Learning the USDMNN windows
The SLGDA for learning the windows of a USDMNN is formalized in Algorithm 2. The initial point $\boldsymbol{W}\in\mathscr{C}$, a batch size $b$, the number $n$ of neighbors to be sampled at each step, and the number of training epochs are fixed. The initial point is stored as the point with minimum validation error visited so far. For each epoch, the validation sample is randomly partitioned in $N/b$ batches, and we denote the validation error of any $\boldsymbol{W}$ on batch $j$ by $L_{v}^{(j)}(\boldsymbol{W})\coloneqq L_{v}^{(j)}(\theta^{\mathbb{A}}_{\boldsymbol{W}})$ which is the empirical error on the $j$-th batch of the validation sample of the USDMNN realized by $\theta^{\mathbb{A}}_{\boldsymbol{W}}$, learned by Algorithm 1.
@@
Algorithm 2 Stochastic lattice gradient descent algorithm for learning the windows of a USDMNN.

### v2

@@ Abstract
There have been attempts to insert mathematical morphology (MM) operators into convolutional neural networks (CNN), and the most successful endeavor to date has been the morphological neural networks (MNN). Although MNN have performed better than CNN in solving some problems, they inherit their black-box nature. Furthermore, in the case of binary images, they are approximations that loose the Boolean lattice structure of MM operators and, thus, it is not possible to represent a specific class of W-operators with desired properties. In a recent work, we proposed the Discrete Morphological Neural Networks (DMNN) for binary image transformation to represent specific classes of W-operators and estimate them via machine learning. We also proposed a stochastic lattice descent algorithm (SLDA) to learn the parameters of Canonical Discrete Morphological Neural Networks (CDMNN), whose architecture is composed only of operators that can be decomposed as the supremum, infimum, and complement of erosions and dilations. In this paper, we propose an algorithm to learn unrestricted sequential DMNN, whose architecture is given by the composition of general W-operators. We illustrate the algorithm in a practical example.
@@
discrete morphological neural networks image processing mathematical morphology U-curve algorithms stochastic lattice descent algorithm
@@ 1 Introduction
Mathematical morphology methods have also been studied in connection with neural networks. The first papers about morphological neural networks (MNN), such as [10, 11, 12, 32], proposed neural network architectures in which the operation performed by each neuron is either an erosion or a dilation. MNN usually have the general structure of neural networks, and their specificity is on the fact that the layers realize morphological operations. Many MNN architectures and training algorithms have been proposed for classification and image processing problems [4, 13, 17, 26, 38]. Special classes of MNN such as morphological/rank neural networks [27, 28], the dendrite MNN [5, 31, 36]; and the modular MNN [3, 37] have been proposed. We refer to [35] and the references therein for a review of the early learning methods based on MNN.
@@
More recently, mathematical morphology methods have been studied in connection with deep neural networks, by either combining convolutional neural networks (CNN) with a morphological operator [21], or by replacing the convolution operations of CNN with basic morphological operators, such as erosions and dilations [16, 18, 25]. Although it has been seen empirically that convolutional layers could be replaced by morphological layers [16], and MNN have shown a better performance than CNN in some tasks [20], they are not more interpretable than an ordinary CNN. In this context, the interpretability is related to the notion of hypothesis space, which is a key component of machine learning theory.
@@
In [22] we proposed the Discrete Morphological Neural Networks (DMNN) for binary image transformation to represent translation invariant and locally defined operators, i.e., W-operators, and estimate them via machine learning. A DMNN architecture is represented by a Morphological Computational Graph that combines operators via composition, infimum, and supremum to build more complex operators. In [22] we proposed a stochastic lattice descent algorithm (SLDA) to train the parameters of Canonical Discrete Morphological Neural Networks (CDMNN) based on a sample of input and output images under the usual machine learning approach. The architecture of a CDMNN is composed only of canonical operations (supremum, infimum, or complement) and operators that can be decomposed as the supremum, infimum, and complement of erosions and dilations with the same structural element. The DMNN is a true mathematical morphology method since it retains the control over the design and the interpretability of results intrinsic to classical mathematical morphology methods, which is a relevant advantage over CNN.
@@
In this paper, we propose an algorithm to train unrestricted sequential DMNN (USDMNN), whose architecture is given by the composition of general W-operators. The algorithm considers the representation of a W-operator by its characteristic Boolean function, which is learned via a SLDA in the Boolean lattice of functions. This SLDA differs from that of [22] which minimizes an empirical error in a lattice of sets and intervals that are the structural elements and intervals representing operators such as erosions, dilations, openings, closings and sup-generating.
@@
Unlike CDMNN, which can be designed from prior information to represent a constrained class of operators, USDMNN are subject to overfitting the data, since it can represent a great class of operators, that may fit the data, but not generalize to new examples. In order to address this issue, we control the complexity of the architecture by selecting from data the window of each W-operator in the sequence (layer). By restricting the window, we create equivalence classes on the characteristic function domain that constrain the class of operators that each layer can represent, decreasing the overall complexity of the operators represented by the architecture and mitigating the risk of overfitting. We propose a SLDA to select the windows by minimizing a validation error in a lattice of sets within the usual model selection framework in machine learning.
@@
In Section 2, we present some notations and definitions, and in Section 3 we formally define the USDMNN that is a particular example of the DMNN proposed by [22]. In Section 4, we present the SLDAs to learn the windows and the characteristic functions of the W-operators in a USDMNN, and in Section 5 we apply the USDMNN to the same dataset used in [22] in order to compare the results. In Section 6, we present the next steps of this research.
@@ 2 W-operators
Let $W$ be a finite subset of $E$. A set operator $\psi$ is called locally defined within a window $W$ if, and only if, $\forall h\in E$, $h\in\psi(X)\iff h\in\psi(X\cap W_{h})$ for $X\in\mathcal{P}(E)$. The collection $\Psi_{W}$ of t.i. operators locally defined within a window $W\in\mathcal{P}(E)$ inherits the complete lattice structure of $(\mathcal{P}(E),\subseteq)$ by setting, $\forall\psi_{1},\psi_{2}\in\Psi_{W}$, {linenomath}
@@
Define by $\Omega=\cup_{W\in\mathcal{P}(E),|W|<\infty}\Psi_{W}$ the collection of all operators from $\mathcal{P}(E)$ to $\mathcal{P}(E)$ that are t.i. and locally defined within some finite window $W\in\mathcal{P}(E)$. The elements of $\Omega$ are called $W$-operators.
@@
$T(\psi)(X)=\begin{cases}1,&\text{ if }o\in\psi(X)\\ 0,&\text{ otherwise}.\end{cases}$ $\psi\in\Psi_{W},X\in\mathcal{P}(W).$ (2)
@@ 3 Unrestricted Sequential Discrete Morphological Neural Networks
In this section, we formally define the USDMNN. In Section 3.1, we define the morphological computational graph of USDMNN. This is a concept introduced in [22] that is related to their architecture, defined in Section 3.2. In Section 3.3, we propose a representation of USDMNN by the windows and characteristic functions of the operators in their layers.
@@ 3.1 Sequential Morphological Computational graph
=$\mathcal{E}=\left\{(\mathfrak{v}_{i},\mathfrak{v}_{1}),(\mathfrak{v}_{n},\mathfrak{v}_{o})\right\}\bigcup\left\{(\mathfrak{v}_{j},\mathfrak{v}_{j+1}):j=1,\dots,n-1\right\},$ (3)
with $\mathcal{C}(\mathfrak{v}_{i})=\mathcal{C}(\mathfrak{v}_{o})=\iota$ and $\mathcal{C}(\mathfrak{v}_{j})\in\Omega,\ j=1,\dots,n$. This computational graph satisfies the axioms of MCG (see [22] for more details).
@@
Denote by $\psi_{\mathcal{G}}(X)$ the output of vertex $\mathfrak{v}_{o}$ when the input of vertex $\mathfrak{v}_{i}$ is $X\in\mathcal{P}(E)$ and let $\psi_{\mathcal{G}}:\mathcal{P}(E)\to\mathcal{P}(E)$ be the set operator generated by MCG $\mathcal{G}$. The operator $\psi_{\mathcal{G}}$ is actually t.i. and locally defined within a window $W_{\mathcal{G}}$ (cf. Proposition 5.1 in [22]). We define the sequential Discrete Morphological Neural Network represented by $\mathcal{G}$ as the translation invariant and locally defined set operator $\psi_{\mathcal{G}}$.
3.2 Sequential Discrete Morphological Neural Networks Architecture
@@ 3.3 Representation of USDMNN
in which $\psi_{j}=\mathcal{C}(\mathfrak{v}_{j}),j=1,\dots,n$. Based on (4), we propose a representation for the class of operators realized by an USDMNN based on the window and characteristic function of the operators $\psi_{j}$.
@@
Since an operator locally defined in $W$ is also locally defined in any $W^{\prime}\supset W$ (cf. Proposition 5.1 in [8]) it follows that $\Theta$ is actually an overparametrization of $\mathcal{H}(\mathcal{A})$ since there are many representations $(W^{\prime},f_{\psi})$ for a same W-operator locally defined in $W\subsetneq F_{d_{i}}$. This overparametrization, discussed more generally in [23], also happens in the representation of CDMNN proposed in [22] and we take advantage of it to propose a SLDA to train USDMNN and mitigate the risk of overfitting.
4 Training USDMNN via the stochastic lattice descent algorithm
@@
The U-curve algorithms [6, 15, 29, 30] were proposed to minimize a U-curve function on a Boolean lattice. In summary, these algorithms perform a greedy search of a Boolean lattice, at each step jumping to the point at distance one with the least value of the function, and stopping when all neighbor points have a function value greater or equal to that of the current point. This greedy search of a lattice, that at each step goes to the direction that minimizes the function, is analogous to the dynamic of the gradient descent algorithm to minimize a function with domain in $\mathbb{R}^{p}$. Inspired by the U-curve algorithms and by the success of stochastic gradient descent algorithms for minimizing overparametrized functions in $\mathbb{R}^{p}$, and following the ideas of [22], we propose a stochastic lattice descent algorithm (SLDA) to train USDMNN. In Figure 1, we present the main ideas of the algorithm, and in Sections 4.1 and 4.2 we formally define it.
Figure 1: Illustration of the deterministic version of the SLDA to (a) learn the USDMNN windows and (b) train a USDMNN with fixed windows. To simplify the illustration, we considered only the possibility of adding a point to a window, and flipping a bit from 0 to 1 of a characteristic function, at each step, even though a point can be erased from a window, and a flipping from 1 to 0 may happen, if they have the least respective error. The training error $L_{t}$ or validation error $L_{v}$ of each point is on top of it.
@@ 4.1 Training a USDMNN with fixed windows
The SLDA for learning the characteristic functions of a USDMNN with fixed windows is formalized in Algorithm 1. The initial point $\theta\in\Theta_{\boldsymbol{W}}$, a batch size $b$, the number $n$ of neighbors to be sampled at each step, and the number of training epochs are fixed. For each epoch, the training sample is randomly partitioned in $N/b$ batches, and we denote the training error on batch $j$ by $L_{t}^{(j)}$. For each batch, $n$ neighbors of $\theta$ are sampled and $\theta$ is updated to the sampled neighbor with the least training error $L_{t}^{(j)}$, that is calculated on the sample batch $j$. Observe that $\theta$ is updated at each batch, so during an epoch, it is updated $N/b$ times. At the end of each epoch, the training error $L_{t}(\theta)$ of $\theta$ on the whole training sample is compared with the error of the point with the least training error visited so far at the end of an epoch, and it is stored as this point if its training error is lesser. After the predetermined number of epochs, the algorithm returns the point with the least training error on the whole sample visited at the end of an epoch.
Observe that Algorithm 1 has two sources of stochasticity: the sampling of neighbors and the sample batches. If $n=n(\theta)\coloneqq|N(\theta)|$ and $b=N$, then this algorithm reduces to the deterministic one illustrated in Figure 1. Furthermore, the complexity of the algorithm is controlled by the number $n$ of sampled neighbors, the batch size $b$ and the number of epochs. See [22] for a further discussion about a SLDA.
Algorithm 1 Stochastic lattice descent algorithm for learning the characteristic functions of a USDMNN with fixed windows $\boldsymbol{W}=\{W_{1},\dots,W_{n}\}$.
@@
4.2 Learning the windows of USDMNN windows
In order to learn the windows of a USDMNN, we apply the SLDA on $\mathscr{C}$, that is a subset of the Boolean lattice $\prod_{i=1}^{n}\{0,1\}^{F_{d_{i}}}$. For each $\boldsymbol{W}\in\mathscr{C}$, let $L_{v}(\boldsymbol{W})\coloneqq L_{v}(\theta^{\mathbb{A}}_{\boldsymbol{W}})$ be the validation error of the USDMNN realized by $\theta^{\mathbb{A}}_{\boldsymbol{W}}$, which was learned by Algorithm 1.
@@ 4.2 Learning the windows of USDMNN windows
The SLDA for learning the windows of a USDMNN is formalized in Algorithm 2. The initial point $\boldsymbol{W}\in\mathscr{C}$, a batch size $b$, the number $n$ of neighbors to be sampled at each step, and the number of training epochs are fixed. For each epoch, the validation sample is randomly partitioned in $N/b$ batches, and we denote the validation error of any $\boldsymbol{W}$ on batch $j$ by $L_{v}^{(j)}(\boldsymbol{W})\coloneqq L_{v}^{(j)}(\theta^{\mathbb{A}}_{\boldsymbol{W}})$ which is the empirical error on the $j$-th batch of the validation sample of the USDMNN realized by $\theta^{\mathbb{A}}_{\boldsymbol{W}}$, learned by Algorithm 1.
@@
Algorithm 2 Stochastic lattice descent algorithm for learning the windows of a USDMNN.